OpenStudy (anonymous):

Okay what am I missing... 3x+8y=24 I isolated the y which I am solving for to the left. 8y=-3x+24. I divided both sides by 8 to get get y by itself. y=-0.5x+3 div by 4. I got y=-0.152+0.75 which = y=0.625. It did not come out correctly. What am I missing. (Must be an integer or fraction.

OpenStudy (anonymous):

3/8 isn't equal 0.5

OpenStudy (anonymous):

why did you divide by 4 after dividing by 8?

OpenStudy (anonymous):

and where did x go?

OpenStudy (anonymous):

I was going off of the example to do it.

OpenStudy (anonymous):

3x+8y = 24 8y = 24-3x y = 3 - (3/8)x

OpenStudy (anonymous):

So once I have that do I plug it in to solve for x to get my full y value?

OpenStudy (anonymous):

what does the question ask for? o,o solve for x? solve for y?

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

y = 3 - (3/8)x yea then this's the answer lol. you can't plug anything in x, unless the question says so

OpenStudy (anonymous):

I was over thinking it... I do that on the easy questions.. always seem to get the hardest ones right and the easy ones wrong. Thanks for the help!